# Multilingual-To-Multimodal (M2M): Unlocking New Languages with Monolingual Text

### v1

@@ Abstract
Multimodal models excel in English, supported by abundant image–text and audio–text data, but performance drops sharply for other languages due to limited multilingual multimodal resources. Existing solutions rely on machine translation, while advances in multilingual text modeling remain underutilized. We introduce M2M, a lightweight alignment method that learns only a few linear layers—using English text alone—to map multilingual text embeddings into multimodal space. Despite its simplicity, M2M matches baseline performance in English (94.9% Recall@10) and achieves strong zero-shot transfer (89.5% Recall@10 averaged across 11 languages, 10 unseen) on XTD Text-to-Image retrieval. Qualitative t-SNE visualizations show that multilingual embeddings align tightly with multimodal representations, while weight analysis reveals that the transformation reshapes embedding geometry rather than performing trivial rotations. Beyond image–text retrieval, M2M generalizes to Audio–Text retrieval and cross-lingual Text-to-Image generation. We release code and checkpoints [^1] GitHub: m2m-codebase/M2M along with multilingual evaluation datasets: MSCOCO Multilingual 30K [^2] HF: piyushsinghpasi/mscoco-multilingual-30k, AudioCaps Multilingual [^3] HF: piyushsinghpasi/audiocaps-multilingual, and Clotho Multilingual [^4] HF: piyushsinghpasi/clotho-multilingual.
@@ 1 Introduction
Humans naturally align information across modalities, associating visual objects with words and sounds. For example, once a person has learned to associate the visual concept of a cat with the English word “cat”, learning that the Spanish word “gato” refers to the same concept allows the object–word association for “gato” to emerge implicitly, without requiring direct visual supervision in Spanish.
@@
In this work, we propose M2M, a simple, data-efficient, and parameter-efficient approach to bridge multilingual text and multimodal latent spaces using English text as a shared anchor. Inspired by how humans learn, our method does not require explicit multimodal signals for each language—English textual data alone is sufficient for alignment. We learn a lightweight projection map, implemented as a few linear layers, while keeping all pretrained encoders frozen and training with MSE and structure-preserving losses. Despite its simplicity, this alignment enables multilingual text representations to participate directly in multimodal tasks, including retrieval and generation, without observing any multilingual image–text or audio–text pairs. Rather than replacing large-scale multilingual multimodal pretraining, our goal is to show that improved latent-space alignment can recover much of the same capability at a fraction of the data and computational cost.
@@ 4 Exploring Alignment Design Space
Figure 3: Effect of scaling train data on XTD eval set for M2M-aligned model- Jina-CLIP-v1 $\times$ M-MPNET.
@@
Figure 4: t-SNE visualization (perplexity = 32) of text embeddings before and after alignment. Marker shapes denote visual clusters and colors indicate languages, with English J-CLIP text embeddings ($z_{e}$ or en_clip) in red. Before alignment (top), text embeddings ($z_{m}$ & $z_{e}$) are fragmented; after alignment (bottom), multilingual captions ($z_{m\to e}$) and J-CLIP text embeddings ($z_{e}$) align closely with shared visual clusters.
@@ 5 Image-Text Retrieval
M2: Jina-CLIP-v1 $\times$ M-MiniLM 86.4 86.7 93.8 88.1 88.4 87.6 80.5 74.8 89 85.2 86.3 90.2 84.9 93.7 57.7 64.5 88.0 85.9
@@
M7: siglip1 $\times$ M-MPNET 86.1 87.2 92.3 86.7 87.2 87.3 80.9 82.8 87.0 83.0 85.6 87.0 85.9 94.5 59.1 65.3 93.5 91.4
@@
Experimental setup. We evaluate English vision–language models ($\mathcal{M}_{e}$): CLIP Radford et al. (2021), Jina-CLIP-v1 Koukounas et al. (2024a), and K-ALIGN Yoon et al. (2022), and multilingual text encoders ($T_{m}$): LaBSE Feng et al. (2020), M-MPNET and M-MiniLM Reimers and Gurevych (2020), and Jina-Text-v3 Sturua et al. (2024). Aligned models are denoted as $\mathcal{M}_{e}\times T_{m}$ (e.g., CLIP $\times$ LaBSE).
We compare M2M-aligned models against: (i) English-only $\mathcal{M}_{e}$ (CLIP, Jina-CLIP-v1, K-ALIGN), and (ii) multilingual multimodal baselines (MMMs): mUSEM3L Aggarwal and Kale (2020), MCLIP-ST (Multilingual CLIP Reimers and Gurevych (2020) from SentenceTransformers [^5] https://www.sbert.net/), MURAL-Large Jain et al. (2021), ALIGN-Base Jia et al. (2021) (via MURAL), XLM-R ViT variants (ViT-L/14, ViT-B/32, ViT-B/16+) & LaBSE ViT-L/14 Carlsson et al. (2022), Jina-CLIP-v2 Koukounas et al. (2024b), and $\text{AltCLIP}_{M9}$ Chen et al. (2022). Supported languages are listed in Appendix C.
@@
Results & Analysis. For the XTD Text-to-Image (T2I) task, the M2M-aligned $\text{Jina-CLIP-v1}\times\text{M-MPNET}$ model (row M4, Table 2) outperforms several MMMs trained on multimodal and/or multilingual paired data (rows T1-T3, T5-T7). For English, our models (row M3 & M4) match English-trained baselines (rows E1-E3). For subsequent comparisons, we use Jina-CLIP-v2 as the SOTA model, as it achieves the best average performance across all languages.
@@ 6 Audio-Text Retrieval
Synthetic Evaluation Datasets. Due to the lack of multilingual audio-text evaluation datasets, we extend the AudioCaps (4,875 captions) and Clotho (5,225 captions) test sets to 33 new languages using machine translation. For 11 Indic languages666bn, gu, hi, kn, ml, mr, ne, pa, ta, te, ur, we use the English-to-Indic translation model from IndicTrans2 Gala et al. (2023), and for 22 other languages777ar, zh-Hans, zh-Hant, cs, nl, fr, de, el, he, id, it, ja, ko, fa, pl, pt, ro, ru, es, tr, uk, vi, we use Aya-23-35B Aryabumi et al. (2024).
@@
Results & Analysis. Table 3 shows that our method generalizes effectively to modalities beyond images. On AudioCaps, our approach performs 6% below the SOTA on Text-to-Audio retrieval (T2A), while on Clotho the gap is 2.3% (T2A).
To understand this gap, we compute Text-to-Text (T2T) Recall@10 on XM3600 (image-text) and AudioCaps (audio-text), leveraging multiple captions per instance. M-MPNET achieves 62.1% T2T Recall@10 on XM3600, comparable to Jina-CLIP-v1 (63.8%), but only 73.8% on AudioCaps, substantially lower than CLAP-General (80.2%). This suggests that M-MPNET, while effective for image-caption encoding, underperforms for audio-caption encoding.
7 Cross-lingual Text-to-Image Generation
@@ Appendix G Curation of Synthetic evaluation dataset
For AYA-23-35B, we use translation prompts to generate synthetic data following Alam et al. (2024). We experiment with zero-shot and 3-shot prompts. We use FLoRes-200 dataset to assess the quality of translation prompts. Zero-shot prompt is fairly straightforward method- we pass the input sentence and prompt the model to generate translation in target language. For 3-shot prompt, for each input english text for which translation has to be generated, we pick 3 examples. These 3 examples are picked from sampling set- created by combining FLoRes-200 validation and test set (excluding current input text). We compute cosine similarity between input text and sampling set using LaBSE, and select top 3 texts and it’s corresponding translation of the target language as a few-shot example. The zero-shot translation prompt performs better on the FLoReS-200 dataset Costa-jussà et al. (2022) across 14 languages111111ar, zho-Hant, fr, de, he, hi, it, jp, ko, pl, ru, es, tr, vi, achieving a mean spBLEU of 39.7 and mean chrF++ of 51.5, compared to the 3-shot prompt with mean spBLEU of 37.2 and mean chrF++ of 47.4. Given these results, we apply the zero-shot prompt to generate Aya-23-35B translations for all 22 languages. Language-wise spBLEU and chrF++ scores for AYA-23-35B are shown in Table 13, and for backtranslated Indic translations are shown in Table 14. Zero-shot prompt and 3-shot prompt templates are listed in Table 11 and 12.

### v2

@@ Abstract
Multimodal models excel in English, supported by abundant image–text and audio–text data, but performance drops sharply for other languages due to limited multilingual multimodal resources. Existing solutions rely on machine translation, while advances in multilingual text modeling remain underutilized. We introduce M2M, a lightweight alignment method that learns only a few linear layers—using English text alone—to map multilingual text embeddings into multimodal space. Despite its simplicity, M2M matches baseline performance in English (94.9% Recall@10) and achieves strong zero-shot transfer (89.5% Recall@10 averaged across 11 languages, 10 unseen) on XTD Text-to-Image retrieval. Qualitative t-SNE visualizations show that multilingual embeddings align tightly with multimodal representations, while weight analysis reveals that the transformation reshapes embedding geometry rather than performing trivial rotations. Beyond image–text retrieval, M2M demonstrates robustness across datasets and tasks, extending to Audio–Text retrieval and Text-to-Image generation. We release code and checkpoints [^1] GitHub: piyushsinghpasi/M2M along with multilingual evaluation datasets: MSCOCO Multilingual 30K [^2] HF: piyushsinghpasi/mscoco-multilingual-30k, AudioCaps Multilingual [^3] HF: piyushsinghpasi/audiocaps-multilingual, and Clotho Multilingual [^4] HF: piyushsinghpasi/clotho-multilingual.
@@ 1 Introduction
Humans naturally align information across modalities, associating visual objects/sounds with words. For example, once a person has learned to associate the visual concept of a cat with the English word “cat”, learning that the Spanish word “gato” refers to the same concept allows the object–word association for “gato” to emerge implicitly, without requiring visual supervision in Spanish.
@@
In this work, we propose M2M, a simple, data-efficient, and parameter-efficient approach to bridge multilingual text and multimodal latent spaces using English text as a shared anchor. Inspired by how humans learn, our method does not require explicit multimodal signals for each language—English textual data alone is sufficient for alignment. We learn a lightweight projection map, implemented as a few linear layers, while keeping all pretrained encoders frozen and training with MSE and structure-preserving losses. Despite its simplicity, this alignment enables multilingual text representations to participate directly in multimodal tasks, including retrieval and generation, without observing any multilingual image–text or audio–text pairs. Rather than aiming to replace large-scale multilingual multimodal pretraining, our goal is to show that improved latent-space alignment can recover much of the same capability at a fraction of the data and computational cost.
@@ 4 Exploring Alignment Design Space
Figure 3: Effect of scaling training data on XTD eval set for M2M-aligned model Jina-CLIP-v1 $\times$ M-MPNET. Recall@10 averaged across all languages in XTD.
@@
Figure 4: t-SNE visualization (perplexity = 32) of text embeddings before and after alignment. Marker shapes denote visual clusters and colors indicate languages, with English Jina-CLIP-v1 text embeddings ($z_{e}$ or en_clip) in red. Before alignment (top), text embeddings ($z_{m}$ & $z_{e}$) are fragmented; after alignment (bottom), multilingual captions ($z_{m\to e}$) and Jina-CLIP-v1 text embeddings ($z_{e}$) align closely with shared visual clusters.
@@
M2: Jina-CLIP-v1 $\times$ M-MiniLM 86.4 86.7 93.8 88.1 88.4 87.6 80.5 74.8 89.0 85.2 86.3 90.2 84.9 93.7 57.7 64.5 88.0 85.9
@@
M7: SigLIP $\times$ M-MPNET 86.1 87.2 92.3 86.7 87.2 87.3 80.9 82.8 87.0 83.0 85.6 87.0 85.9 94.5 59.1 65.3 93.5 91.4
@@
5 Image-Text Retrieval
Experimental setup. We evaluate English vision–language models ($\mathcal{M}_{e}$): CLIP Radford et al. (2021), Jina-CLIP-v1 Koukounas et al. (2024a), K-ALIGN Yoon et al. (2022), & SigLIP [^5] HF: google/siglip-so400m-patch14-384 Zhai et al. (2023), and multilingual text encoders ($T_{m}$): LaBSE Feng et al. (2020), M-MPNET and M-MiniLM Reimers and Gurevych (2020), and Jina-Text-v3 Sturua et al. (2024). Aligned models are denoted as $\mathcal{M}_{e}\times T_{m}$ (e.g., CLIP $\times$ LaBSE).
We compare M2M-aligned models against: (i) English-only $\mathcal{M}_{e}$ (CLIP, Jina-CLIP-v1, K-ALIGN), and (ii) multilingual multimodal baselines (MMMs): mUSEM3L Aggarwal and Kale (2020), MCLIP-ST (Multilingual CLIP Reimers and Gurevych (2020) from SentenceTransformers [^6] https://www.sbert.net/), MURAL-Large Jain et al. (2021), ALIGN-Base Jia et al. (2021) (via MURAL), XLM-R ViT variants (ViT-L/14, ViT-B/32, ViT-B/16+) & LaBSE ViT-L/14 Carlsson et al. (2022), Jina-CLIP-v2 Koukounas et al. (2024b), $\text{AltCLIP}_{M9}$ Chen et al. (2022), SigLIP, SigLIP2777HF: google/siglip2-so400m-patch14-384 Tschannen et al. (2025), and LLM2CLIP [^8] HF: microsoft/LLM2CLIP-Openai-L-14-336 Huang et al. (2024). Supported languages are listed in Appendix C.
@@ 5 Image-Text Retrieval
Results & Analysis. For the XTD Text-to-Image (T2I) task, the M2M-aligned $\text{Jina-CLIP-v1}\times\text{M-MPNET}$ model (row M4, Table 2) outperforms several multilingual multimodal models trained with multilingual and/or multimodal paired data (rows T1–T3, T5–T7, T11). On English, our aligned models (rows M3 and M4) match the performance of English-trained baselines (rows E1–E3), indicating that alignment does not degrade monolingual performance. Importantly, these gains arise solely from representation alignment, as no new multilingual or multimodal supervision is introduced during training. For subsequent comparisons, we adopt Jina-CLIP-v2 as the state-of-the-art reference model, since it achieves the strongest average performance across different languages and datasets.
@@ 6 Audio-Text Retrieval
Synthetic Evaluation Datasets. Due to the lack of multilingual audio–text evaluation datasets, we extend the AudioCaps and Clotho test sets to 33 additional languages using machine translation. In particular, we translate 4,875 AudioCaps captions (although evaluation was performed on only 4,785 captions because 18 audio files were unavailable during dataset creation) and 5,225 Clotho captions. For 11 Indic languages999bn, gu, hi, kn, ml, mr, ne, pa, ta, te, ur, we use the English-to-Indic translation model from IndicTrans2 Gala et al. (2023). For the remaining 22 languages101010ar, zh-Hans, zh-Hant, cs, nl, fr, de, el, he, id, it, ja, ko, fa, pl, pt, ro, ru, es, tr, uk, vi, we use Aya-23-35B model Aryabumi et al. (2024).
@@
Results & Analysis. Table 3 shows that our method generalizes effectively to modalities beyond images. On AudioCaps, our approach performs 6% below the SOTA on Text-to-Audio retrieval (T2A), while on Clotho the gap is 2.3% (T2A). To understand this gap, we compute Text-to-Text (T2T) Recall@10 on XM3600 (image-text) and AudioCaps (audio-text), leveraging multiple captions per instance. M-MPNET achieves 62.1% T2T Recall@10 on XM3600, comparable to Jina-CLIP-v1 (63.8%), but only 73.8% on AudioCaps, substantially lower than CLAP-General (80.2%). This suggests that M-MPNET, while effective for image-caption encoding, underperforms for audio-caption encoding.
@@ 8 Conclusion
Acknowledgements
The authors thank Prof. Preethi Jyothi (Indian Institute of Technology, Bombay) for her continued support and for fostering the research environment that enabled this research. Her guidance and encouragement were instrumental throughout the course of this research. We are grateful to Pavan Kalyan for extensive discussions, careful proofreading, and thoughtful feedback on early drafts of the paper, which greatly improved the clarity, structure, and presentation of this work, and to Isha Pandey for her help in reviewing, refining, and strengthening the rebuttal responses.
@@ Appendix G Curation of Synthetic evaluation dataset
For AYA-23-35B, we use translation prompts to generate synthetic data following Alam et al. (2024). We experiment with zero-shot and 3-shot prompts. We use FLoRes-200 dataset to assess the quality of translation prompts. Zero-shot prompt is fairly straightforward method- we pass the input sentence and prompt the model to generate translation in target language. For 3-shot prompt, for each input english text for which translation has to be generated, we pick 3 examples. These 3 examples are picked from sampling set- created by combining FLoRes-200 validation and test set (excluding current input text). We compute cosine similarity between input text and sampling set using LaBSE, and select top 3 texts and it’s corresponding translation of the target language as a few-shot example. The zero-shot translation prompt performs better on the FLoReS-200 dataset Costa-jussà et al. (2022) across 14 languages141414ar, zho-Hant, fr, de, he, hi, it, jp, ko, pl, ru, es, tr, vi, achieving a mean spBLEU of 39.7 and mean chrF++ of 51.5, compared to the 3-shot prompt with mean spBLEU of 37.2 and mean chrF++ of 47.4. Given these results, we apply the zero-shot prompt to generate Aya-23-35B translations for all 22 languages. Language-wise spBLEU and chrF++ scores for AYA-23-35B are shown in Table 13, and for backtranslated Indic translations are shown in Table 14. Zero-shot prompt and 3-shot prompt templates are listed in Table 11 and 12.